Determine the value of k so that the line passing through the points (k,5) and (4,-1) is perpendicular to the line containing points (2,-2) and (1,-5).
Express your result as an exact number, using a fraction if necessary.

Answer by mananth(16946)

(k,5) and (4,-1) is perpendicular to the line containing points (2,-2) and (1,-5).
(k,5) and (4,-1)
slope = (y2-y1)/(x2-x1)
plug the values
(-1-5)/(4-k)= -6/(4-k)
(2,-2) and (1,-5).
slope = (-5-(-2)/(1-2)=-3/-1 = 3

The slopes of perpendicular lines are negative reciprocal of each other
m1*m2 = -1

((-6)/(4-k)) *3 =-1
-18= -1(4-k)
-18 = k-4
k=-14


=